OpenStudy (anonymous):

2x2 + 7x + 6 = 0 How do I found out if this is: double Root Real and Rational Root Real and Irrational Root Imaginary Root

jimthompson5910 (jim_thompson5910):

From 2x^2+7x+6 we can see that a=2, b=7, and c=6 D=b^2-4ac D=(7)^2-4(2)(6) D=49-4(2)(6) D=49-48 D=1 So the discriminant is D=1 Since the discriminant is greater than zero, this means that there are two real solutions. Also, because the discriminant is a perfect square, this means that you'll have two rational solutions.